Will is w years old
Ben is 3 years older.
Jan is twice as old as Ben.
If you add togather the ages of Will, Ben, and Jan the total comes to 41 years.
Write an equation and solve it to work out how old Will, Ben, and Jan are.

Answer by ankor@dixie-net.com(22740)

Will is w years old
Ben is 3 years older.
Jan is twice as old as Ben.
If you add together the ages of Will, Ben, and Jan the total comes to 41 years.
Write an equation and solve it to work out how old Will, Ben, and Jan are.
:
b = w + 3
j = 2b
therefore
j = 2(w+3)
j = 2w + 6
:
If you add together the ages of Will, Ben, and Jan the total comes to 41 years.
w + b + j = 41
Replace b with (w+3), replace j = (2w+6)
w + w + 3 + 2w + 6 = 41
4w + 9 = 41
4w = 41 - 9
4w = 32
w = 32/4
w = 8 yrs is will
you can find the ages of Ben and jan, check the total